Hi please answer asap Thank you !

Gator Beach Marts, a chain of convenience grocery stores in the Fort Lauderdale area, has store hours that fluctuate from monReq 1A Req 1B Req ЗА Req 3B Req 4 What is the variable utility cost per hour of operation? Variable cost per hourReq 1A Req 1B Req ЗА Req 3B Req 4 Construct an Excel spreadsheet and use the Excel commands to perform a least-squares regresWhat is the variable utility cost per hour of operation? (Do not round intermediate calculations. Round the intercept to theDuring July, the store will be open 300 hours. Predict the stores total utility cost for July using each of the cost-estimat
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Gator Beach
High Low method
Variable cost= (Highest activity cost- Lowest activity cost)/ (Highest activity units- Lowest activity units)
Answer 1 Month Hours of operation Utility costs
Highest activity March                        800.00                  3,900.00
Lowest activity June                        300.00                  1,900.00
Change                        500.00                  2,000.00
Amount $
Variable cost per hour                       4.00
Fixed cost: Highest activity cost- (Highest units * Variable cost per unit)
Fixed cost: 3900-(800*4)
Fixed cost                  700.00
Answer 1 a
Y= 700+ 4.x
Answer 1 b
Variable cost per hour = $ 4.
Answer 4 Amount $
Hours in July                  300.00
Variable cost per hour                       4.00
Variable cost               1,200.00
Fixed cost                  700.00
Total cost               1,900.00
Answer 3 a
Regression method
The equation of a line is y=mx+b
X= Hours
y=total cost
m= (N Σ(xy) − Σx Σy)/ (N Σ(x2) − (Σx)2)
Month Hours (x) Costs (y) x2 xy
January                  450.00                     2,520.00              202,500.00    1,134,000.00
February                  600.00                     3,100.00              360,000.00    1,860,000.00
March                  800.00                     3,900.00              640,000.00    3,120,000.00
April                  400.00                     2,400.00              160,000.00       960,000.00
May                  350.00                     2,050.00              122,500.00       717,500.00
June                  300.00                     1,900.00                90,000.00       570,000.00
Total               2,900.00                  15,870.00          1,575,000.00 8,361,500.00
N= number of months= 6.
m= (N Σ(xy) − Σx Σy)
(N Σ(x2) − (Σx)2)
m= 6*8361500- 2900*15870
6*1575000- 2900^2
m= 50169000- 46023000
9450000-8410000
m=       4,146,000.00
      1,040,000.00
m=                       3.99
b= (Σy − m Σx)/N
b= 15870- 3.99*2900
                      6.00
b= 15870- 11571
                      6.00
b=               4,299.00
                      6.00
b=                  717.00
So fixed cost per month is $ 717.
Answer 3 b
Variable cost per hour is $ 3.99.
Answer 4 Amount $
Hours in July                  300.00
Variable cost per hour                       3.99
Variable cost               1,197.00
Fixed cost                  717.00
Total cost               1,914.00
Final Answer 4 Amount $
High Low method               1,900.00
Regression method               1,914.00